# HD 44892: The youngest (or oldest?) gas-harbouring debris disc around an intermediate mass star

### v1

@@ Abstract
We present the first detections of gas around a 2.1 Myr old debris disc-bearing intermediate-mass star, HD 44892. Through our analysis we deem the alternative age of 800 Myr unlikely. Gas is detected both in 12CO (2-1) emission through ALMA Band 6 observations and absorption in Ca ii K and H seen with high-resolution UVES spectroscopy. The star exhibits a 12 m fractional excess of , placing it in the transition stage between protoplanetary and debris discs, at a level comparable to HD 141569. Our detection of 1.3 mm emission yields the dust mass of 0.019 0.009 M⊕ assuming 115 K temperature. Assuming a gas temperature of 20 K, the resulting CO gas mass is (7.86 2.05) 10-5 . For a warmer gas of 50 K it increases to (1.62 0.17) 10-4 . From the 13CO and C18O non-detections we derive upper limits which suggest this CO gas mass estimate should be treated as a lower limit, because the CO emission might be optically thick. From the analysis of UVES spectroscopic data we find variability in the spectrum of HD 44892, in both calcium Ca ii K and H lines, which is absent from sodium Na i D1 and D2 lines, and can be interpreted by transiting circumstellar gas. Through comparison to spectra of nearby stars we exclude the possibility of these features being caused by interstellar absorption. Both the dust mass, which is within an order of magnitude of HD 141569, and the gas mass derived here indicate a late gas dispersal stage of the protoplanetary disc.
@@
stars: pre-main-sequence – planets and satellites: formation – circumstellar matter1 Introduction
@@
While the dust content of debris discs has been extensively studied, less is known about the presence of gas. Numerous papers have searched for and examined this gas presence (e.g., Rebollido et al., 2022; Moór et al., 2017; Cataldi et al., 2023; Hales et al., 2022; Marino et al., 2016; Moór et al., 2019; Iglesias et al., 2018); however, many questions remain unanswered. Perhaps the biggest question of all is: where does this gas come from? Two main scenarios have been proposed. This gas is either primordial, leftover from the parent protoplanetary disc, or secondary, produced alongside dust, through processes such as planetesimal collisions, volatile sublimation, and photodesorption (Hughes et al., 2018). Recent findings regarding the shielding effects in discs suggest that CO gas may accumulate in relatively high amounts from secondary means alone (Kral et al., 2019). Nonetheless, the primordial origin has not yet been disproved, especially as the effects of gas mixing complicate the case for the secondary gas scenario, making shielding less efficient (Marino et al., 2022), and models of Nakatani et al. (2021) suggest that, in case of A-type stars, the gas can survive as a protoplanetary remnant.
@@
As of December 2024, 45 debris discs containing gas have been discovered either through gas emission or absorption. Over 20 of them have been observed with ALMA to search for CO and other gas emission, a full list of which can be found in Table 2. Interestingly, the majority of all gas detections in debris discs have been around A-type stars (see Figure 1). For most of these systems, secondary origin models can readily explain the observed low CO gas masses. However, a few systems, such as 49 Cet (45 Myr, Moór et al., 2019; Pearce et al., 2022), HD21997 (42 Myr, Kóspál et al., 2013; Moór et al., 2013), HD141569 (5 Myr, Miley et al., 2018), and HD32297 (< 30 Myr, Moór et al., 2019), which are intermediate-mass stars, exhibit significantly higher CO gas masses, which the current secondary models are not able to fully explain. In these cases, the origin of the gas is proposed to be primordial or a mixture of primordial and secondary (Cataldi et al., 2023). Secondary origin is most certain around older systems, such as the 440 Myr Fomalhaut (Matrà et al., 2017a) or the 1-2 Gyr Crv (Marino et al., 2017), where primordial gas cannot survive over such long timescales. Pic, despite its younger age of 20 Myr, is another example of a system with secondary gas, in which gas produced in planetesimal collisions and the destruction of volatile bodies has been observed (Dent et al., 2014; Matrà et al., 2017b).
@@
The primordial origin has been suggested for discs where the gas is present in substantial amounts - quantities that are not yet fully explainable by models of secondary origin - and where the parent star is very young (primordial gas is possible in systems < 50 Myr-old in models of Nakatani et al., 2021), hence the gas might not have yet been fully cleared from the system (Smirnov-Pinchukov et al., 2022). Such is the case for HD 141569 (Miley et al., 2018), a 5 Myr star harbouring a ‘hybrid’ disc, name used for gas-bearing discs with protoplanetary disc levels of likely primordial gas, but debris-disc levels of secondary dust. These discs are similar to dusty debris discs but retain some primordial gas from the earlier evolutionary stage, hence are hybrid between protoplanetary and debris discs (Péricaud et al., 2017). The dissipation of primordial gas is still ongoing in these systems, placing them just at the end of protoplanetary disc dispersal. In the case of HD 141569, the dust is of secondary origin or mixed, with a relatively low dust mass and IR excess for a protoplanetary disc but high for a debris disc (Figure 2), while the gas is suspected to be primordial and at a mass comparable to the least-massive protoplanetary discs (PPDs). HD 141569 is thus the only known gas-harbouring debris disc found around a pre-main-sequence (pre-MS) intermediate-mass star (IMS) with a measured CO mass to date.
@@ 2 HD 44892
In this paper, we present ALMA observations of HD 44892 (also known as HIP 30414), an A9/F0IV type star (SIMBAD, Wenger et al., 2000) located 191 1 pc away (Gaia Collaboration et al., 2023, 2016). HD 44892 has a temperature of 7500 500K, a luminosity of 60.0 20.0 , stellar mass of 2.87 , and an age of 2.1 Myr (Iglesias et al., 2023). The pre-MS star HD 44892 has undergone the transition from convective to radiative structure at 1 Myr. Gontcharov (2006) have calculated the radial velocity of the star to be 20.9 0.6 km s-1, which is the value adopted in this study. A more recent estimate yields a value of 20.88 1.85 km s-1 (D. Iglesias, priv. comm.), which is consistent with previous findings. There are no mentions in the literature suggesting HD 44892 is part of a binary or multiple star system; it is also a field star (Gagné et al., 2018).
Ishihara et al. (2017) identified it as a warm debris disc candidate based on its IR excess at 18 m found through spectral energy distribution (SED) fitting, and it was first identified as a hybrid disc candidate in Iglesias et al. (2023). HD 44892 has a fractional IR excess at 12 m of 7.86 (Iglesias et al., 2023), placing it between debris () and protoplanetary () discs (Wyatt et al., 2015; Iglesias et al., 2023), where also a number of gas-bearing debris discs have been found (Figure 2, right).
There is a disagreement in the literature regarding the stellar parameters of HD 44892, particularly about its age. Its location on the HR diagram can be fit by either pre- or post-MS tracks. Past studies, such as Zorec & Royer (2012) and Bochanski et al. (2018), have classified it as a post-MS object (570-580 Myr with a mass of 2.50-2.60 ). Iglesias et al. (2023) present the fit to the pre-MS tracks yielding the 2.1 Myr and an alternative fit to the post-MS tracks with 800 Myr, with a mass of 2.29 . They conclude that the possibility of the older age is less likely. In this paper, we adopt the pre-MS age of 2.1 Myr and present arguments against the alternative age of 800 Myr in later sections.
@@ 3.1 ALMA observations and data reduction
We processed all of the data using CASA 6.5.4-9 (McMullin et al., 2007) and CARTA 4.0.0 (Comrie et al., 2021). Continuum emission of HD 44892 was imaged using the task ‘tclean’ with Briggs weighting and a robust parameter set to 0.3, which we found to provide the highest signal-to-noise ratio and source emission over the range of -2 to 2 explored. Our target’s emission fits entirely in the region where primary beam is maximum, hence no primary beam correction was applied. No self-calibration was applied to the data as the source is too weak in continuum to provide reliable phase or amplitude solutions.
For line data, we used the CASA task ‘uvcontsub’ to subtract the continuum from the line emission data. Using the ‘plotms’ task (amplitude vs channel), we identified SPWs and frequencies at which line emission was potentially present. We imaged both SPWs using the ‘tclean’ task with Briggs weighting and a robust parameter set to 0.3, which produced the highest signal-to-noise ratio for the 12CO detection. We did not detect any emission in the individual channels of SPW containing 13CO and C18O.
@@ 3.2.1 Continuum
For continuum of HD 44892 our data reduction resulted in a synthesised beam size of 044 (84 au) 038 (72 au) at a position angle (PA) of . We detected continuum emission of 0.136 0.015 mJy (9 ), shown in Figure 3. The continuum flux was measured only from inside the 3 contour. Using task ‘imfit’, which is part of the CASA package, we found that the continuum emission of HD 44892 is marginally resolved, as the major axis of the fitted Gaussian component (058 008, 110 au) extends beyond that of the beam (045, 84 au, measured in the same direction). This allows us to examine the structure in the continuum emission, which seems to deviate from circular, being more extended and asymmetric (Figure 3).
@@
where is the full width at half maximum (FWHM) of the synthesized beam, and SNR is the signal-to-noise ratio of the peak target flux. For our data we found = 45 miliarcseconds. The apparent offset of the continuum emission from the stellar position is 34 miliarcseconds, which is less than the calculated , suggesting that it is most likely due to the limited astrometric accuracy rather than physical displacement.
@@ 3.2.2 Gas
For 12CO emission, the resulting synthesized beam size is 046 (88 au) 040 (76 au), at a PA of -80.6. Figure 4 shows the spectrum of the detected gas emission, which has the typical double-peaked profile associated with gas rotating in a Keplerian disc. Fitting a double-peaked Gaussian curve to this spectrum reveals a midpoint at km s-1. The first peak is at -5.45 km s-1 and the second at 6.75 km s-1. This x-axis of this plot has been corrected for the adopted heliocentric velocity of the system, 20.90 km s-1 and the star’s radial velocity in the LSR frame of .
@@
From moment 0 (Figure 5, integrated intensity made using the CASA task ‘immoments’ within range -8.49-15.33 km s-1) we found an integrated line flux of 96.92 16.22 mJy km s-1 (6.0 ). Therefore, we decided to re-image the data, this time binning three channels together during tclean (nchan = 3). This reduced the noise in our image and made the emission more prominent. The moment 0 noise decreased slightly to 14.10 mJy km s-1, and the integrated line flux to 93.06 mJy km s-1, improving the signal-to-noise (6.6 ).
@@
We then also used the Keplerian Mask tool222https://github.com/richteague/keplerian_mask, which produces a rotating mask based on the system’s parameters that a Keplerian disc should follow. To find the parameters needed, we used the ‘imfit’ task to obtain deconvolved Gaussian parameters, resulting in an inclination of 56.1 19.9 and PA of 76.2 63.4. The deconvolved disk size was found to be 034 010 for the major axis and 019 008 for the minor axis, hence the 12CO emission is spatially unresolved. The stellar mass range was adopted from Iglesias et al. (2023) as 2.87 . Using these values as initial constraints, we explored combinations of the parameters (within their uncertainty ranges) to best match the 12CO emission of HD 44892. The parameters that resulted in the best fit were an inclination of 40, PA of 17.2, and stellar mass of 3.27 . The 12CO emission follows this rotation, indicating that the emission lines up with theoretical positions and confirming it is from a disc rotating around the parent star (Figure 6). Using the Keplerian Mask during cleaning has improved the signal-to-noise once more, in moment 0 this resulted in an integrated line flux () of 95.47 13.49 mJy km s-1 (7.1 , Figure 5). Here we used channels where gas emission was most prominent, corresponding to the frequency range of 230.5262 GHz (15.33 km s-1) to 230.5445 GHz (-8.49 km s-1).
@@
Knowing the channels (velocities) at which 12CO emission is present, we created a moment 0 map for SPW containing 13CO and C18O. The emission from these isotopes would allow for a more precise CO gas mass calculation, as they are likely optically thin, ensuring the 12CO emission traces molecular density, rather than being convoluted with temperature effects. Since no emission was visible in the individual channels of this spectral window, we used the same frequency range as for 12CO. However, this has resulted in a non-detection of both 13CO and C18O.
@@
Figure 5 shows the full extent of 12CO gas emission (contour levels) overlaid on the velocity distribution of the disc. The visible velocity gradient, from red- to blue-shifted, indicates gas rotation. This, along with the disc being located around the star’s position, confirms the gas is rotating around the star and the emission is not from a different source. The gas emission is relatively symmetric, with the highest intensity near the centre, gradually decreasing toward the edges.
@@ 4.1 UVES observations and data reduction
We obtained high-resolution observations of HD 44892 using UVES (Ultraviolet and Visual Echelle Spectrograph; Dekker et al. 2000) at the VLT (Very Large Telescope), on 13th and 27th September 2023 (PI Iglesias, ID 112.262P.001). VLT is part of the European Southern Observatory and is located in Cerro Paranal in Chile. Data was taken in both Blue and Red arms, centered at 390 nm and 580 nm, respectively. We used the 05 for these observations to achieve high resolution (R74000 for the Red arm and R65000 for the Blue arm). HD 44892 was observed each time for 1 minute 48 seconds. Data of one of the comparison stars used, HIP 28877, is part of the same observing programme. The observations were reduced using the UVES EsoReflex pipeline (Ballester et al., 2000; Freudling et al., 2013) version 6.4.1.
Additional spectra from the ESO archive333http://archive.eso.org were obtained for two other comparison stars (HD 43544, HD 46035), using high-resolution spectrographs HARPS (Mayor et al., 2003) and FEROS Kaufer et al. (1999), also on the VLT. These were science-ready products from the ESO archive and reduced with each instrument standard pipeline. We used Molecfit (Smette et al., 2015; Kausch et al., 2015) to correct for telluric contamination as the region around the Na i doublet is highly affected by the Earth’s atmosphere. Barycentric radial velocity corrections were also applied to the spectra.

### v2

@@ Abstract
We present the first detections of gas around a 2.1 Myr old debris disc-bearing intermediate-mass star, HD 44892. Gas is detected both in 12CO (2-1) emission through ALMA Band 6 observations and in absorption in Ca ii K and H, seen with high-resolution UVES spectroscopy. 13CO and C18O (2-1) are not detected. The star exhibits a 12 m fractional excess of , placing it in the transition stage between protoplanetary and debris discs. Our detection of 1.3 mm emission yields the dust mass of 0.019 0.009 M⊕ assuming 115 K temperature. The 13CO non-detection places an upper limit on CO gas mass of 10-2 . The 12CO detection yields a CO gas mass of (7.86 2.05) 10-5 or (1.62 0.17) 10-4 assuming a gas temperature of 20 K or 50 K, respectively. These should be considered as lower limits since 12CO emission may be optically thick. With UVES, we find variability in Ca ii K and H lines, which can be interpreted by transiting circumstellar gas, ruling out interstellar absorption as their origin. Both the dust mass, which is within an order of magnitude of HD 141569, and the gas mass derived here indicate a late gas dispersal stage of the protoplanetary disc. Through our analysis we deem the alternative age of 800 Myr unlikely.
@@
protoplanetary discs – submillimetre: planetary systems – stars: pre-main-sequence – circumstellar matter – planets and satellites: formation1 Introduction
@@
While the dust content of debris discs has been extensively studied, less is known about the presence of gas. Numerous papers have searched for and examined this gas presence (e.g., Rebollido2022The36546; Moor2017MolecularStars; Cataldi2023PrimordialALMA; Hales2022ALMADisk; Marino2016ExocometaryRing; Moor2019New32297; Iglesias2018DebrisOrigin); however, many questions remain unanswered. Perhaps the biggest question of all is: where does this gas come from? Two main scenarios have been proposed. This gas is either primordial, leftover from the parent protoplanetary disc, or secondary, produced alongside dust, through processes such as planetesimal collisions, volatile sublimation, and photodesorption (Hughes2018DebrisVariability). Recent findings regarding the shielding effects in discs suggest that CO gas may accumulate in relatively high amounts from secondary means alone (Kral2019ImagingDiscs). Nonetheless, the primordial origin has not yet been disproved, especially as the effects of gas mixing complicate the case for the secondary gas scenario, making shielding less efficient (Marino2022VerticalLifetime), and models of Nakatani2021PhotoevaporationRemnants suggest that, in case of A-type stars, the gas can survive as a protoplanetary remnant.
@@
As of August 2025, 45 debris discs containing gas have been discovered either through gas emission or absorption. Over 20 of them have been observed with ALMA to search for CO and other gas emission, a full list of which can be found in Table 2. Interestingly, the majority of all gas detections in debris discs have been around A-type stars (see Figure 1). For most of these systems, secondary origin models can readily explain the observed low CO gas masses. However, a few systems, such as 49 Cet (45 Myr, Moor2019New32297; Pearce2022PlanetDiscs), HD 21997 (42 Myr, Kospal2013ALMA21997; Moor2013ALMA21997), HD 141569 (5 Myr, Miley2018Unlocking141569), and HD 32297 (< 30 Myr, Moor2019New32297), which are intermediate-mass stars, exhibit significantly higher CO gas masses, which the current secondary models are not able to fully explain. In these cases, the origin of the gas is proposed to be primordial or a mixture of primordial and secondary (Cataldi2023PrimordialALMA). Secondary origin is most certain around older systems, such as the 440 Myr Fomalhaut (Matra2017ExocometaryObservations) or the 1-2 Gyr Crv (Marino2017ALMAPlanets), where primordial gas cannot survive over such long timescales. Pic, despite its younger age of 20 Myr, is another example of a system with secondary gas, in which gas produced in planetesimal collisions and the destruction of volatile bodies has been observed (Dent2014MolecularDisk; Matra2017DetectionComets).
@@
The primordial origin has been suggested for discs where the gas is present in substantial amounts - quantities that are not yet fully explainable by models of secondary origin - and where the parent star is very young (primordial gas is possible in systems < 50 Myr-old in models of Nakatani2021PhotoevaporationRemnants), hence the gas might not have yet been fully cleared from the system (Smirnov-Pinchukov2022LackGas). Such is the case for HD 141569 (Miley2018Unlocking141569), a 5 Myr star harbouring a ‘hybrid’ disc, name used for gas-bearing discs with protoplanetary disc levels of likely primordial gas, but debris-disc levels of secondary dust. These discs are similar to dusty debris discs but retain some primordial gas from the earlier evolutionary stage, hence are hybrid between protoplanetary and debris discs (Pericaud2017TheDisks). The dissipation of primordial gas is still ongoing in these systems, placing them just at the end of protoplanetary disc dispersal. In the case of HD 141569, the dust is of secondary origin or mixed, with a relatively low dust mass and IR excess for a protoplanetary disc but high for a debris disc (Figure 2), while the gas is suspected to be primordial and at a mass comparable to the least-massive protoplanetary discs (PPDs). HD 141569 is thus the only known gas-harbouring debris disc found around a pre-main-sequence (pre-MS) intermediate-mass star (IMS) with a measured CO mass to date.
@@ 2 HD 44892
In this paper, we present ALMA observations of HD 44892 (also known as HIP 30414), an A9/F0IV type star (SIMBAD, Wenger2000TheDatabase) located 191 1 pc away (Vallenari2023iGaia/i3; Prusti2016TheMission). HD 44892 has a temperature of 7500 500K, a luminosity of 60.0 20.0 , stellar mass of 2.87 , and an age of 2.1 Myr (Iglesias2023X-shooterEvolution). The pre-MS star HD 44892 has undergone the transition from convective to radiative structure at 1 Myr. Gontcharov2006PulkovoSystem have calculated the radial velocity of the star to be 20.9 0.6 km s-1, which is the value adopted in this study. A more recent estimate yields a value of 20.88 1.85 km s-1 (D. Iglesias, priv. comm.), which is consistent with previous findings. There are no mentions in the literature suggesting HD 44892 is part of a binary or multiple star system; it is also a field star (Gagne2018BANYAN.Pc).
Ishihara2017FaintIRSF identified it as a warm debris disc candidate based on its IR excess at 18 m found through spectral energy distribution (SED) fitting, and it was first identified as a hybrid disc candidate in Iglesias2023X-shooterEvolution. HD 44892 has a fractional IR excess at 12 m of 7.86 (Iglesias2023X-shooterEvolution), placing it between debris () and protoplanetary () discs (Wyatt2015FiveDisk; Iglesias2023X-shooterEvolution), where also a number of gas-bearing debris discs have been found (Figure 2, right).
There is a disagreement in the literature regarding the stellar parameters of HD 44892, particularly about its age. Its location on the HR diagram can be fit by either pre- or post-MS tracks. Past studies, such as Zorec2012RotationalStars and Bochanski2018FundamentalGaia, have classified it as a post-MS object (570-580 Myr with a mass of 2.50-2.60 ). Iglesias2023X-shooterEvolution present the fit to the pre-MS tracks yielding the 2.1 Myr and an alternative fit to the post-MS tracks with 800 Myr, with a mass of 2.29 . They conclude that the possibility of the older age is less likely. In this paper, we adopt the pre-MS age of 2.1 Myr and present arguments against the alternative age of 800 Myr in later sections.
@@ 3.1 ALMA observations and data reduction
We processed all of the data using CASA 6.5.4-9 (McMullin2007CASAApplications) and CARTA 4.0.0 (Comrie2018CARTA:Astronomy). Continuum emission of HD 44892 was imaged using the task ‘tclean’ with Briggs weighting and a robust parameter set to 0.3, which we found to provide the highest signal-to-noise ratio and source emission over the range of -2 to 2 explored. Our target’s emission fits entirely in the region where primary beam is maximum, hence no primary beam correction was applied. No self-calibration was applied to the data as the source is too weak in continuum to provide reliable phase or amplitude solutions.
For line data, we used the CASA task ‘uvcontsub’ to subtract the continuum from the line emission data. Using the ‘plotms’ task (amplitude vs channel), we identified SPWs and frequencies at which line emission was potentially present. We imaged both SPWs using the ‘tclean’ task with Briggs weighting and a robust parameter set to 0.3, which produced the highest signal-to-noise ratio for the 12CO detection. We did not detect any emission in the individual channels of the SPW containing 13CO and C18O.
@@ 3.2.1 Continuum
For continuum of HD 44892 our data reduction resulted in a synthesised beam size of 044 (84 au) 038 (72 au) at a position angle (PA) of . We detected continuum emission of 0.136 0.016 mJy (9 ), shown in Figure 3. The continuum flux was measured only from inside the 3 contour. Using task ‘imfit’, which is part of the CASA package, we found that the continuum emission of HD 44892 is marginally resolved, as the fitted Gaussian component has a FWHM of major axis of 058 008, 110 au, which extends beyond that of the beam (043, 82 au, measured in the same direction). The drop in amplitude at longest UV distances (Figure 4), confirms that the most extended emission from the source is being resolved.
@@
where is the full width at half maximum (FWHM) of the synthesized beam, and SNR is the signal-to-noise ratio of the peak target flux. For our data we found = 45 milliarcseconds. The apparent offset of the continuum emission from the stellar position is 34 milliarcseconds, which is less than the calculated , suggesting that it is most likely due to the limited astrometric accuracy rather than physical displacement.
@@ 3.2.2 Gas
For 12CO emission, the resulting synthesized beam size is 046 (88 au) 040 (76 au), at a PA of -80.6. Figure 5 shows the spectrum of the detected gas emission, which has the typical double-peaked profile associated with gas rotating in a Keplerian disc. Fitting a double-peaked Gaussian curve to this spectrum reveals a midpoint at km s-1. The first peak is at -5.45 km s-1 and the second at 6.75 km s-1. This x-axis of this plot has been corrected for the star’s radial velocity in the LSR frame of .
@@
From moment 0 (Figure 6, integrated intensity made using the CASA task ‘immoments’ within range -8.49-15.33 km s-1) we found an integrated line flux of 96.92 16.22 mJy km s-1 (6.0 ). Therefore, we decided to re-image the data, this time binning three channels together during tclean (nchan = 3). This reduced the noise in our image and made the emission more prominent. The moment 0 map was made using the same velocity range as in the non-binned case, resulting in a slightly lower noise of 14.10 mJy km s-1, and the integrated line flux of 93.06 mJy km s-1, improving the signal-to-noise (6.6 ).
@@
We then also used the Keplerian Mask tool222https://github.com/richteague/keplerian_mask, which produces a rotating mask based on the system’s parameters that a Keplerian disc should follow. To search the parameter space to be used in the tool, we are guided by the aspect ratio of the deconvolved disc size. We used the ‘imfit’ task and obtained the deconvolved Gaussian parameters: 034 010 for the major axis and 019 008 for the minor axis (PA of 76.2 63.4). This suggests an inclination of 56.1 19.9. We therefore varied the inclination from 36 to 76 in the Keplerian Mask tool. The stellar mass range was adopted from Iglesias2023X-shooterEvolution as 2.87 and Keplerian rotation computed accordingly in the tool. Using these values as initial constraints, we explored combinations of the parameters (within their uncertainty ranges) to best match the 12CO emission of HD 44892. The parameters that resulted in the best fit were an inclination of 40, PA of 17.2, stellar mass of 3.27 , and an outer radius of 04 (76 au). The 12CO emission follows this rotation, indicating that the emission lines up with theoretical positions and confirming it is from a disc rotating around the parent star (Figure 7). Using the Keplerian Mask during cleaning has improved the signal-to-noise once more, in moment 0 this resulted in an integrated line flux () of 95.47 13.49 mJy km s-1 (7.1 , Figure 6). Here we used channels where gas emission was most prominent, corresponding to the frequency range of 230.5262 GHz (15.33 km s-1) to 230.5445 GHz (-8.49 km s-1).
@@
Knowing the channels (velocities) at which 12CO emission is present, we created a moment 0 map for the SPW containing 13CO and C18O. The emission from these isotopes would allow for a more precise CO gas mass calculation, as they are likely optically thin, ensuring the 12CO emission traces molecular density, rather than being convoluted with temperature effects. Since no emission was visible in the individual channels of this spectral window, we used the same velocity range as for 12CO. However, this has resulted in a non-detection of both 13CO and C18O.
@@
Figure 6 shows the full extent of 12CO gas emission (contour levels) overlaid on the velocity distribution of the disc. The visible velocity gradient, from red- to blue-shifted, indicates gas rotation. This, along with the disc being located around the star’s position, confirms the gas is rotating around the star and the emission is not from a different source. The gas emission appears compact and follows the beam shape, with the highest intensity near the centre, as expected for a spatially unresolved source.
@@ 4.1 UVES observations and data reduction
We obtained high-resolution observations of HD 44892 using UVES (Ultraviolet and Visual Echelle Spectrograph; Dekker2000DesignObservatory) at the VLT (Very Large Telescope), on 13th and 27th September 2023 (PI Iglesias, ID 112.262P.001). VLT is part of the European Southern Observatory and is located in Cerro Paranal in Chile. Data was taken in both Blue and Red arms, centred at 390 nm and 580 nm, respectively. We used the 05 for these observations to achieve high resolution (R74000 for the Red arm and R65000 for the Blue arm). HD 44892 was observed each time for 1 minute 48 seconds. Data of one of the comparison stars used, HIP 28877, is part of the same observing programme. The observations were reduced using the UVES EsoReflex pipeline (Ballester2000ThePipeline; Freudling2013AutomatedAstronomy) version 6.4.1.
Additional spectra from the ESO archive333http://archive.eso.org were obtained for two other comparison stars (HD 43544, HD 46035), using high-resolution spectrographs HARPS (Mayor2003SettingHARPS) and FEROS Kaufer1999CommissioningLa-Silla. at the La Silla Observatory. These were science-ready products from the ESO archive and reduced with each instrument standard pipeline. We used Molecfit (Smette2015Molecfit:Correction; Kausch2015Molecfit:Correction) to correct for telluric contamination as the region around the Na i doublet is highly affected by the Earth’s atmosphere. Barycentric radial velocity corrections were also applied to the spectra.